### v1

# On the required number of electrodes for uniqueness and convex reformulation in an inverse coefficient problem

Abstract
We introduce a computer-assisted proof for uniqueness and global reconstruction for the inverse Robin transmission problem, where the corrosion function on the boundary of an interior object is to be determined from current-voltage measurements on the boundary of an outer domain. We consider the shunt electrode model where, in contrast to the standard Neumann boundary condition, the applied electrical current is only partially known. The aim is to determine the corrosion coefficient with a finite number of measurements.
In this paper, we present a numerically verifiable criterion that ensures unique solvability of the inverse problem, given a desired resolution. This allows us to explicitly determine the required number and position of the required electrodes. Furthermore, we will present an error estimate for noisy data. By rewriting the problem as a convex optimisation problem, our aim is to develop a globally convergent reconstruction algorithm. The results extend the recent research in [1] to the shunt electrode model and establish a more precise criterion for unique solvability.
  • October 2024
Keywords: Inverse problems, shunt electrode model, global convergence
1 Introduction
In our study, we focus on non-destructive impedance-based corrosion detection that holds the challenge of reconstructing an unknown Robin transmission coefficient on a known interior boundary from current-voltage measurements taken at electrodes attached to the outer boundary.
In this work, we aim to prove uniqueness and global reconstruction in a non-linear inverse elliptic coefficient problem. To the best knowledge of the authors, this work presents the first method for explicitly calculating the required number of electrodes for a resolution guarantee in a non-linear spatially dependent coefficient reconstruction problem.
Practitioners usually attempt to solve such reconstruction problems with a regularized data fitting approach. The non-convexity of the residual function presents a significant challenge for nonlinear inverse problems, which is why, in general, only a few global reconstruction algorithms are known. In [1, 2], a novel approach to overcome non-convexity and achieve global convergence is presented for the idealized model, where standard Neumann boundary conditions are considered. Our study translates these results to the shunt electrode model, which extends the standard Neumann boundary problem to a realistic electrode model, where the applied electrical current is only partially known. We provide a criterion to check whether the number of electrodes is sufficient for unique solvability of the inverse problem and provide an equivalent reformulation by rewriting the problem as a convex semidefinite program. We are able to explicitly determine the required number of electrodes for the reconstruction guarantee of the unknown Robin parameter.
By evaluating a finite number of forward solutions, and their derivatives, the criterion can be explicitly and numerically verified for a given desired resolution, so the proof for uniqueness and global reconstruction is computer-assisted. Additionally, the criterion provides explicit error estimates for noisy data. Finally, we genrealize the criterion for unique solvability in [1] to sharper assumptions.
The Calderón problem cf. [3], is the problem of reconstructing information about the interior of a domain from measurements at its boudary. Traditionally, a prominent application of the Calderón problem is electrical impedance tomography (EIT), see eg. [4] for an overview. In EIT, one seeks to reconstruct information from a finite number of current-voltage measurements. There are recent theoretical achievements in global reconstruction of the unknown conductivity parameter with infinitely many measurements cf. [5, 6]. However, there are only very few achievements in global reconstruction approaches with finitely many measurements cf. [7, 8, 9, 10] and [11] for the fractional case. The Robin problem, that is considered in [1], is a special case where one seeks to reconstruct a parameter on a small subdomain. For the Robin case theoretical results, such as Runge approximation and the technique of localized potentials, hold in general eg. cf. [12]. Theoretical investigations of uniqueness for the Robin case have mainly focused on the idealized infinite-dimensional setting, where the unknown coefficient function is to be determined with infinite resolution from infinitely many measurements as in [12]. The works [1, 2] address the practically relevant scenario of a finite resolution and finitely many measurements, where a measurement corresponds to measuring the exact Dirichlet data from applied Neumann boundary values. In [12], the authors have established Lipschitz stability results for cases involving finitely many unknowns and infinitely many measurements.
In this work, we consider the Robin problem with a realistic electrode model, where electrodes are attached to the outer boundary of the domain (see [13]). There are different types of electrode models, most prominently the complete electrode model. We consider a slightly simplified model, the shunt model, where the impedance between the electrodes and the body is neglected. In [14, 15, 16, 17], some theoretical fundations have been layed. In [9, 18, 19, 20] the authors aim to approximate the idealized model with electrode models and hereby translate the theoretical results to the electrode model case. However, in EIT little is known about characterizing the required number of electrodes to achieve a desired resolution.
Figure 1: The domain $\Omega=\Omega_{1}\cup\Omega_{2}$.
2 Main results
Let $\Omega\subset\mathbb{R}^{d}$ be an open, bounded domain with Lipschitz boundary $\partial\Omega$, $\Omega_{1}$ be an open subset of $\Omega$ with Lipschitz boundary $\Gamma:=\partial\Omega_{1}\subset\Omega$, and $\Omega_{2}:=\Omega\setminus\Omega_{1}$ is connected, see Figure 1. Furthermore, let $\mathcal{E}_{1},\ldots\mathcal{E}_{m}$ be measurable subsets of $\partial\Omega$ with nonzero measure.
For a piecewise-constant conductivity $\sigma=\sigma_{1}\chi_{\Omega_{1}}+\sigma_{2}\chi_{\Omega_{2}}$ with given $\sigma_{1},\sigma_{2}>0$ and a Robin transmission coefficient $\gamma\in L^{\infty}_{+}(\Gamma)$ we consider the following problem with applied currents $I_{k}\in\mathbb{R}$ for $k=1,\ldots,m$
$\left\{\begin{array}[]{rcll}-\nabla\cdot\left(\sigma\nabla u\right)&=&0&\ \ \textrm{on}\ \ \Omega_{1}\cup\Omega_{2},\\ \int_{\mathcal{E}_{k}}\sigma\partial_{\nu}u\,\mathrm{d}s&=&I_{k}&\ \ \textrm{for}\ \ k=1,\ldots,m,\\ \sigma\partial_{\nu}u&=&0&\ \ \textrm{on}\ \ \partial\Omega\setminus\bigcup_{k=1,\ldots,m}\mathcal{E}_{k},\\ \left\llbracket u\right\rrbracket&=&0&\ \ \textrm{on}\ \ \Gamma,\\ \left\llbracket\sigma\partial_{\nu}u\right\rrbracket&=&\gamma u&\ \ \textrm{on}\ \ \Gamma,\\ u\big{|}_{\mathcal{E}_{k}}&=&\textrm{const.}=:U_{k}&\ \ \textrm{for}\ \ k=1,\ldots,m,\end{array}\right.$ (1)
where $\nu$ is the unit normal vector on $\Gamma$ or $\partial\Omega$ pointing outward of $\Omega_{1}$ or $\Omega$,
$\llbracket\varphi\rrbracket:=\textrm{trace}\left(\varphi\big{|}_{\Omega_{2}}\right)-\textrm{trace}\left(\varphi\big{|}_{\Omega_{1}}\right)$and $\llbracket\sigma\partial_{\nu}\varphi\rrbracket_{\Gamma}=\partial_{\nu}\left(\sigma_{2}\varphi\big{|}_{\Omega_{2}}\right)-\partial_{\nu}\left(\sigma_{1}\varphi\big{|}_{\Omega_{1}}\right)$.
A known current $I_{k}$ is applied to the electrode $\mathcal{E}_{k}$ for $k=1,...,m$ and the necessary voltages $U_{k}$ are then measured on the same electrodes. This is called the shunt model. It corresponds, in contrast to the complete electrode model, to the idealistic case of perfect conduction between the body and the electrodes. The variational form of the shunt electrode model is given by
$b(u,v):=\int_{\Omega}\nabla u\nabla v\ {\mathrm{d}}x+\int_{\Gamma}\gamma u\leavevmode\nobreak\ v\ {\mathrm{d}}s=\sum_{k=1}^{m}I_{k}\left(v\big{|}_{\mathcal{E}_{k}}\right)\qquad\textrm{for all}\leavevmode\nobreak\ v\in H_{\square}^{1}(\Omega),$
with $H_{\square}^{1}(\Omega)=\{v\in H^{1}(\Omega):\leavevmode\nobreak\ v\big{|}_{\mathcal{E}_{k}}=\textrm{const.}\quad\textrm{for}\leavevmode\nobreak\ k=1,\ldots,m\}$. Since $b$ is continuous and coercive and $H_{\square}^{1}(\Omega)$ is a closed subspace of $H^{1}(\Omega)$ the non-linear partial differential Equation 1 is uniquely solvable for all $\gamma\in L^{\infty}_{+}(\Gamma)$ by Lax–Milgram Theorem (set forth in [14]). For a piecewise constant coefficient function $\gamma\in L^{\infty}_{+}(\Gamma)$ on a partition $\Gamma_{1},\ldots,\Gamma_{n}$, $n\geq 2$, where
$\Gamma_{1},\ldots,\Gamma_{n},\leavevmode\nobreak\ \textrm{are pairwise disjoint, measurable and}\leavevmode\nobreak\ \bigcup\limits_{k=1}^{n}\Gamma_{k}=\Gamma,$
one can define a forward map $\mathcal{F}$ to model the measurements. :+$\mathcal{F}\colon\mathbb{R}^{n}_{+}\to\mathbb{S}^{m}_{+}$ is given by
[…]
Here $\mathbb{S}^{m}_{+}$ denotes the space of the symmetric positive definite matrices and we identify the vector $\gamma\in\mathbb{R}^{n}_{+}$ with the piecewise constant function $\left(\sum\limits_{k=1}^{n}\gamma_{k}\chi_{\Gamma_{k}}\right)\in L^{\infty}_{+}(\Gamma)$. Note that $\mathcal{F}(\gamma)\in\mathbb{S}^{m}_{+}$ since
$I^{T}\mathcal{F}(\gamma)J=b(u_{\gamma}^{(I)},u_{\gamma}^{(J)})\qquad\textrm{for all}\leavevmode\nobreak\ I,J\in\mathbb{R}^{m}$
and $b$ is symmetric and coercive.
The physical interpretation of $\mathcal{F}(\gamma)$ is the measurement of the resulting voltage $U\in\mathbb{R}^{m}$ along the electrodes corresponding to the applied currents $I\in\mathbb{R}^{m}$. In particular the $(jk)$-th component of $\mathcal{F}(\gamma)$ is the voltage at $\mathcal{E}_{j}$ when a current is applied to $\mathcal{E}_{k}$. So we will refer to $\mathcal{F}$ as the measurement operator and to $\mathcal{F}(\gamma)$ as the current-voltage measurements. The resulting inverse problem states as:
Reconstruct the corrosion $\textit{Reconstruct the corrosion parameter}\leavevmode\nobreak\ \gamma\leavevmode\nobreak\ \textit{from the measurements}\leavevmode\nobreak\ \mathcal{F}(\hat{\gamma}).$ (3)
Even if one can prove unique solvability of the inverse problem 3, major difficulties happen to be ill-posedness and non-convexity of the natural data fitting approach. In Figure 2 we apply the generic MATLAB solver lsqnonlin to minimize the least squared residual function
$\|\mathcal{F}(\gamma)-\mathcal{F}(\hat{\gamma})\|_{F}^{2}$
to demonstrate the difficulty of global reconstruction algorithms.
Figure 2: Logaritmic error $\log_{10}(\|\gamma^{(N)}-\hat{\gamma}\|)$ of the standard data fitting for different initial values $\gamma^{0}=(\gamma_{1},\gamma_{2})$, with $\hat{\gamma}=(100,100)$.
After fixing known a priori bounds $a<\gamma_{i}<b$ for $i=1,\ldots,n$ we will be able to present a criterion that guarantees for a fixed number of electrodes $m$ that:
  • The inverse problem 3 is uniquely solvable.
  • Provide an explicit Lipschitz stability constant.
  • There is a globally convergent reconstruction algorithm.
The basic idea to provide such a criterion is to use monotonicity relations as in [1].
Let us define the required inequalities. In the following for $x,y\in\mathbb{R}^{n}$ the relation $x\leq y$ is understood pointwise, i.e. if $x_{i}\leq y_{i}$ for $i=1,\ldots,n$, and for $A,B\in\mathbb{S}^{m}$ the relation $A\preceq B$ is understood in the Loewner sense, i.e. if $\left(B-A\right)$ is positive semidefinite. The two main results state as follows:
Theorem 1.
For a resolution of dimension $n\in\mathbb{N}$, $m\in\mathbb{N}$ fixed electrodes and $n,m\geq 2$ let :+$\mathcal{F}\colon\mathbb{R}^{n}_{+}\to\mathbb{S}_{m}$ be the measurement operator $\mathcal{F}(\gamma)I=U$ as in 2. Furthermore let $a,b\in\mathbb{R}$ with $0<a<b$. If
$\mathcal{F}^{\prime}\left(z_{j,k}\right)d_{j}\npreceq 0\quad\textrm{for all}\quad k\in\{2,\ldots,K\},\leavevmode\nobreak\ j\in\{1,\ldots,n\},$ (4)
where
$z_{j,k}:=\frac{a}{2}e_{j}^{\prime}+\left(a+k\frac{a}{4}\right)e_{j}\in\mathbb{R}^{n}_{+},\leavevmode\nobreak\ d_{j}:=\frac{2b-a}{a}e_{j}^{\prime}-\frac{1}{2}e_{j}\in\mathbb{R}^{n}$
and $K:=\max\left(\lceil\frac{4b}{a}\rceil-3,2\right)$, then the following holds:
  • (a)
    $\mathcal{F}(\gamma)^{\prime}\in\mathcal{L}\left(\mathbb{R}^{n},\mathbb{S}_{m}\right)$ is injective for all $\gamma\in[a,b]^{n}$ and for all $0\neq d\in\mathbb{R}^{n}$,
    $\frac{\|\mathcal{F}(\gamma)^{\prime}d\|_{2}}{\|d\|_{\infty}}\geq\lambda:=\min_{\mathop{j=1,\ldots,n}\limits_{k=2,\ldots,K}}\lambda_{\max}\left(\mathcal{F}^{\prime}(z_{j,k})d_{j}\right)>0.$
  • (b)
    :+$\mathcal{F}\colon\mathbb{R}^{n}_{+}\to\mathbb{S}^{m}$ is injective and
    $\|\mathcal{F}(\gamma_{1})-\mathcal{F}(\gamma_{2})\|_{2}\geq\lambda\|\gamma_{1}-\gamma_{2}\|_{\infty}$
    for all $\gamma_{1},\gamma_{2}\in[a,b]^{n}$ i.e. the inverse problem
    determine $\hat{\gamma}\in[a,b]^{n}$ from the knowledge of $\mathcal{F}(\hat{\gamma})$
    is uniquely solvable.
Criterion 4 allows us to decide whether a number of electrodes is sufficient for unique solvability of the inverse problem 3. The evaluation of the criterion already provides a Lipschitz satbility constant of the inverse problem. The results in theorem 1 generalize the uniqueness result in Theorem 1 in [1] to tighter assumptions. For the proof of Theorem 1 in Section 2.1 and Section 2.2, we will generalize Lemma 1 and Lemma 2 in [1].
For a fixed number of electrodes $m$ and a resolution of dimension $n$, one can explicitly check criterion 4, and therefor decide whether the number and position of the electrodes is sufficient for unique solvability and Lipschitz stability of the inverse problem. Our second result provides a criterion for a convex reformulation of the considered inverse problem and gives an error estimate for noisy data. Note that the assumptions in the following theorem imply the assumptions of Theorem 1.
Theorem 2.
Let $n,m\geq 2$, :+$\mathcal{F}\colon\mathbb{R}^{n}_{+}\to\mathbb{S}_{m}$ be the measurement operator $\mathcal{F}(\gamma)I=U$ and $0<a<b$. For the weaker assumptions
$z_{j,k}:=\frac{a}{2}e_{j}^{\prime}+\left(a+k\frac{a}{4(n-1)}\right)e_{j}\in\mathbb{R}^{n}_{+},\leavevmode\nobreak\ d_{j}:=\frac{2b-a}{a}(n-1)e_{j}^{\prime}-\frac{1}{2}e_{j}\in\mathbb{R}^{n}$
and $K:=\max\left(\lceil\frac{4(n-1)b}{a}\rceil-4n-3,2\right)$ it holds:
If
$\mathcal{F}^{\prime}\left(z_{j,k}\right)d_{j}\\ \npreceq 0\quad\textrm{for all}\quad k\in\{2,\ldots,K\},\leavevmode\nobreak\ j\in\{1,\ldots,n\},$ (5)
then the following holds additionaly to the asserations of Theorem 1:
  • (a)
    The inverse problem
    determine $\hat{\gamma}\in[a,b]^{n}$ from the knowledge of $\mathcal{F}(\hat{\gamma})$
    is uniquely solvable and $\hat{\gamma}$ is the unique minimizer of the convex optimization problem:
    $\texttt{minimize}\leavevmode\nobreak\ \sum_{i=1}^{n}|\gamma_{i}|\leavevmode\nobreak\ \textrm{subject to}\leavevmode\nobreak\ \gamma\in[a,b]^{n},\leavevmode\nobreak\ \mathcal{F}(\gamma)\preceq\mathcal{F}(\hat{\gamma}).$ (6)
  • (b)
    For $\hat{\gamma}\in[a,b]^{n}$, $\delta>0$ and $Y^{\delta}\in\mathbb{S}^{m}$, with $\|\mathcal{F}(\hat{\gamma})-Y^{\delta}\|_{2}\leq\delta$ the convex problem:
    $\texttt{minimize}\leavevmode\nobreak\ \sum_{i=1}^{n}|\gamma_{i}|\leavevmode\nobreak\ \textrm{subject to}\leavevmode\nobreak\ \gamma\in[a,b]^{n},\leavevmode\nobreak\ \mathcal{F}(\gamma)\preceq Y^{\delta}+\delta I$
    possesses a minimum, and every such minimum $\gamma^{\delta}$ fulfills
    $\|\hat{\gamma}-\gamma^{\delta}\|_{\infty}\leq\frac{2\delta(n-1)}{\lambda},$
    with $\lambda:=\min_{j,k}\lambda_{\max}\left(\mathcal{F}^{\prime}(z_{j,k})d_{k}\right)$.
Note that Theorem 2 follows immediately from [1, Theorem 1] with monotonicity and convexity relations proven in Lemma 1. The fundamental concept of the proof in [1] revolves around a converse monotonicity result. It is shown that if $\mathcal{F}$ satisfies condition 5, then
$\mathcal{F}(y)\preceq\mathcal{F}(x)\quad\textrm{implies}\quad\sum_{i=1}^{n}(y_{j}-x_{j})>0\qquad\textrm{for all}\leavevmode\nobreak\ x,y\in\mathbb{R}^{n}_{+},\leavevmode\nobreak\ x\neq y.$
This provides a reformulation of the inverse problem 3 as a convex semi-definite optimization problem. But here the constant $C=(n-1)$ which extends the condition of Theorem 1 is substantial.
Lemma 1.
The forward map :+$\mathcal{F}\colon\mathbb{R}^{n}_{+}\to\mathbb{S}^{m}$ is infinitely many times differentiable and :+$\mathcal{F}^{\prime}\colon\mathbb{R}^{n}_{+}\to\mathcal{L}\left(\mathbb{R}^{n},\mathbb{S}^{m}\right)$ is given by
$I^{T}\left(\mathcal{F}^{\prime}(\gamma)\delta\right)J=-\sum_{i=1}^{n}\delta_{i}\int_{\Gamma_{i}}u_{\gamma}^{(I)}u_{\gamma}^{(J)}\ {\mathrm{d}}x.$
Moreover $\mathcal{F}$ is monotonically non-increasing and convex, i.e. it holds
$\begin{array}[]{rcll}\mathcal{F}^{\prime}(\gamma)\delta&\preceq&0&\quad\textrm{for all }\gamma\in\mathbb{R}^{n}_{+},\,0\leq\delta\in\mathbb{R}^{n},\\[8.0pt] \mathcal{F}(\gamma)-\mathcal{F}(\gamma^{(0)})&\succeq&\mathcal{F}^{\prime}(\gamma^{(0)})\left(\gamma-\gamma^{(0)}\right)&\quad\textrm{for all }\gamma,\gamma^{(0)}\in\mathbb{R}^{n}_{+}.\end{array}$
Proof.
This follows directly from Lemma 2 and Corollary 1 in [21]. ∎
Monotonicity and convexity can also be written as
$0<\gamma^{(0)}\leq\gamma^{(1)}\qquad\textrm{implies}\qquad\mathcal{F}\left(\gamma^{(0)}\right)\succeq\mathcal{F}\left(\gamma^{(1)}\right),$
and for all $\gamma^{(0)},\gamma^{(1)}\in\mathbb{R}^{n}_{+}$, $t\in[0,1]$
$\mathcal{F}\left((1-t)\gamma^{(0)}+t\gamma^{(1)}\right)\preceq(1-t)\mathcal{F}\left(\gamma^{(0)}\right)+t\mathcal{F}\left(\gamma^{(1)}\right).$
The equivalence to the differential characterization in Lemma 1 is shown in [21, Lemma 2].
2.1 A criterion for uniqueness and Lipschitz sability
To prove Theorem 1 we first derive a sufficient criterion for unique solvability of the inverse problem that utilizes directional derivatives in arbitary points $x\in[a,b]^{n}$.
Lemma 2.
Let $m,n\in\mathbb{N}$, $n,m\geq 2$ and $F\colon\mathbb{R}^{n}_{+}\to\mathbb{S}_{m}$ be continuously differentiable, monotonically non-increasing and convex.
  • (a)
    If for some $x\in\mathbb{R}^{n}_{+}$
    $F^{\prime}\left(x\right)\left(e_{j}^{\prime}-e_{j}\right)\npreceq 0\quad\textrm{for all}\quad j\in\{1,\ldots,n\},$
    then $F(x)^{\prime}\in\mathcal{L}\left(\mathbb{R}^{n},\mathbb{S}_{m}\right)$ is injective for all $x\in[a,b]^{n}$. Moreover, for all $0\neq d\in\mathbb{R}^{n}$,
    $\frac{\|F(x)^{\prime}d\|_{2}}{\|d\|_{\infty}}\geq\lambda_{x}:=\min_{j=1,\ldots,n}\lambda_{\max}\left(F^{\prime}\left(x\right)\left(e_{j}^{\prime}-e_{j}\right)\right)>0.$ (7)
  • (b)
    If $x,y\in[a,b]^{n}$ fulfill 7, then
    $\|F(x)-F(y)\|_{2}\geq\lambda\|x-y\|_{\infty}$
    with $\lambda:=\min\{\lambda_{x},\lambda_{y}\}>0$.
Proof.
This proof is copletely analogous to [1, Lemma 1]. But since it is short and simple, we state it for the sake of completeness.
  1. (a)
    Let $d\in\mathbb{R}^{n}$ with $\|d\|_{\infty}=1$. Then at least one of the entries of $d$ must be either $1$ or $-1$, so that there exists $j\in\{1,\ldots,n\}$ with either
    $d\leq e_{j}^{\prime}-e_{j}\quad\textrm{or}\quad-d\leq e_{j}^{\prime}-e_{j}.$
    Since $F$ is monotonically non-increasing we obtain
    $F^{\prime}\left(x\right)d\succeq F^{\prime}\left(x\right)\left(e_{j}^{\prime}-e_{j}\right)\quad\textrm{or}\quad F^{\prime}\left(x\right)(-d)\succeq F^{\prime}\left(x\right)\left(e_{j}^{\prime}-e_{j}\right)$
    for every $x\in[a,b]^{n}$ so that $F^{\prime}\left(x\right)d$ has at least one positive eigenvalue that is larger than $\lambda_{\max}\left(F^{\prime}\left(x\right)\left(e_{j}^{\prime}-e_{j}\right)\right)$ or smaller than $-\lambda_{\textrm{max}}\left(F^{\prime}\left(x\right)\left(e_{j}^{\prime}-e_{j}\right)\right)$ for every $j=1,\ldots,n$. This proves $(a)$.
  2. (b)
    With the same argument as in $(a)$ we have for $x\neq y$ either
    $\frac{x-y}{\|x-y\|_{\infty}}\leq e_{j}^{\prime}-e_{j}\quad\textrm{or}\quad\frac{y-x}{\|x-y\|_{\infty}}\leq e_{j}^{\prime}-e_{j}.$
    Hence by convexity and anti-monotonicity we have that either
    $\begin{array}[]{rl}\displaystyle\frac{F(x)-F(y)}{\|x-y\|_{\infty}}&\succeq F^{\prime}\left(y\right)\frac{x-y}{\|x-y\|_{\infty}}\succeq F^{\prime}\left(y\right)\left(e_{j}^{\prime}-e_{j}\right)\quad\textrm{or}\\[12.0pt] \displaystyle\frac{F(y)-F(x)}{\|x-y\|_{\infty}}&\succeq F^{\prime}\left(x\right)\frac{y-x}{\|x-y\|_{\infty}}\succeq F^{\prime}\left(x\right)\left(e_{j}^{\prime}-e_{j}\right),\end{array}$
    which shows that
    $\|F(x)-F(y)\|_{2}\geq\lambda_{y}\|x-y\|_{\infty}\quad\textrm{or}\quad\|F(x)-F(y)\|_{2}\geq\lambda_{x}\|x-y\|_{\infty},$
    so that (b) is proven.
For inverse coefficient problems with finitely many measurements, the assumption
$\mathcal{F}^{\prime}\left(x\right)\left(e_{j}^{\prime}-e_{j}\right)\npreceq 0$ (8)
can be interpreted in the sense of localized potentials, since
$I^{T}\mathcal{F}^{\prime}\left(x\right)\left(e_{j}^{\prime}-e_{j}\right)I=\int_{\Gamma_{i}}\left(u_{\gamma}^{(I)}\right)^{2}\ {\mathrm{d}}x-\sum_{j=1\atop j\neq i}\int_{\Gamma_{j}}\left(u_{\gamma}^{(I)}\right)^{2}\ {\mathrm{d}}x$
for all $I\in\mathbb{R}^{m}$. This expression is positive if one can ensure that the solution is large on some part of $\Gamma$ and small on the other parts. For some applications such as the idealized Robin problem ([1, Section 3]) it follows from [1, Theorem 2(b)] that these assumptions hold for all $\gamma\in[a,b]^{n}$ when enough measurements are being used, i.e. if $m$ is suffenciently large. For the shunt model 1 it remains to be shown that potentials can be localized with a sufficient number of electrodes, such that the assumptions of Theorem 1 are fulfilled. In [18] the authors provide approximation results of the continuum model by the complete electrode model. These should hold analogously for the shunt electrode model in the Robin case.
2.2 Checking the assumption with only finitely many tests
To practically determine how many measurements are enough, we will now rewrite the assumption so that they require evaluations of $\mathcal{F}^{\prime}$ for only finitely many points. We finalize the proof of Theorem 1 by generalizing Lemma 2 in [1]. Therefore we argue that if one checks the localized potential property 8 for finitely many points, then the assertions of Lemma 2 directly follow for the whole required domain. The key idea is that we can replace the evaluation points of $F^{\prime}$ by using the convexity property
$F^{\prime}(x)(y-x)\succeq F(y)-F(x)\succeq F^{\prime}(y)(y-x)$
and the direction of the derivative by the monotonicity property
$F^{\prime}(x)d_{2}\succeq F^{\prime}(x)d_{1}\qquad\textrm{for}\leavevmode\nobreak\ d_{1}\geq d_{2}.$
Lemma 3.
Let $F:\mathbb{R}_{+}^{n}\rightarrow\mathbb{S}_{m},n,m\geq 2$, be continuously differentiable, convex and monotonically non-increasing, $b\geq a>0$, $C>0$, $d_{j}:=\frac{2b-a}{a}Ce_{j}^{\prime}-\frac{1}{2}e_{j}\in\mathbb{R}^{n}$. Then, for all $j\in\{1,\ldots,n\}$ and $x\in[a,b]^{n}$, there exists $t\in$ $\left[a+\frac{a}{2C},b+\frac{a}{2C}\right]\subset\mathbb{R}$, so that for all $0\leq\delta\leq\frac{a}{4C}$,
$F^{\prime}(x)\left(Ce_{j}^{\prime}-e_{j}\right)\succeq F^{\prime}\left(\frac{a}{2}e_{j}^{\prime}+(t-\delta)e_{j}\right)d_{j}.$
Proof.
Let $j\in\{1,\ldots,n\}$ and $x\in[a,b]^{n}$. We define $t:=x_{j}+\frac{a}{2c}$. Then for all $0\leq\delta\leq\frac{a}{4c}$
$\begin{array}[]{rl}Ce_{j}^{\prime}-e_{j}&=\displaystyle\frac{2C}{a}\left(\frac{a}{2}e_{j}^{\prime}+(x_{j}-t)e_{j}\right)\leq\frac{2C}{a}\left(x-\frac{a}{2}e_{j}^{\prime}-te_{j}\right)\\[10.0pt] &\leq\displaystyle\frac{2C}{a}\left(x-\left(\frac{a}{2}e_{j}^{\prime}+(t-\delta)e_{j}\right)\right)\end{array}$
and
$\begin{array}[]{rl}&\displaystyle\frac{2C}{a}\left(x-\left(\frac{a}{2}e_{j}^{\prime}+(t-\delta)e_{j}\right)\right)\leq\displaystyle\frac{2C}{a}\left(\left(b-\frac{a}{2}\right)e_{j}^{\prime}+\left(x_{j}-t+\delta\right)e_{j}\right)\\[10.0pt] &\quad=\displaystyle\frac{2b-a}{a}(n-1)e_{j}^{\prime}+\frac{2(n-1)}{a}\left(-\frac{a}{2C}+\delta\right)e_{j}\\ &\quad\leq\displaystyle\frac{2b-a}{a}Ce_{j}^{\prime}-\frac{1}{2}e_{j}=d_{j}.\end{array}$
so that we obtain from monotonicity and convexity
$\begin{array}[]{rl}F^{\prime}(x)&\left(Ce_{j}^{\prime}-e_{j}\right)\succeq\displaystyle\frac{2C}{a}F^{\prime}(x)\left(x-\left(\frac{a}{2}e_{j}^{\prime}+(t-\delta)e_{j}\right)\right)\\[5.0pt] &\succeq\displaystyle\frac{2C}{a}\left(F(x)-F\left(\frac{a}{2}e_{j}^{\prime}+(t-\delta)e_{j}\right)\right)\\[10.0pt] &\succeq\displaystyle\frac{2C}{a}F^{\prime}\left(\frac{a}{2}e_{j}^{\prime}+(t-\delta)e_{j}\right)\left(x-\left(\frac{a}{2}e_{j}^{\prime}+(t-\delta)e_{j}\right)\right)\\[10.0pt] &\succeq F^{\prime}\left(\frac{a}{2}e_{j}^{\prime}+(t-\delta)e_{j}\right)d_{j}.\end{array}$
Corollary 1.
Let $C>0$ and $j\in\{1,\ldots,n\}$. Choose $K\geq 2$ so that $a+K\frac{a}{4C}\geq b+\frac{a}{4C}$. If
$\left(F^{\prime}\left(z_{j,k}\right)\left(d_{j}\right)\right)\npreceq 0\qquad\textrm{for all}\leavevmode\nobreak\ k\in\{2,\ldots,K\},$
$z_{j,k}=\frac{a}{2}e_{j}^{\prime}+\left(a+k\frac{a}{4C}\right)e_{j}\in\mathbb{R}^{n}_{+},\leavevmode\nobreak\ d_{j}=\frac{2b-a}{a}Ce_{j}^{\prime}-\frac{1}{2}e_{j}\in\mathbb{R}^{n}$, then
$F^{\prime}(x)\left(Ce_{j}^{\prime}-e_{j}\right)\npreceq 0\qquad\textrm{for all}\leavevmode\nobreak\ x\in[a,b]^{n}.$
Proof.
Since $a+K\frac{a}{4C}\geq b+\frac{a}{4C}$, we have for every $t\in\left[a+\frac{a}{2C},b+\frac{a}{2C}\right]$ that $\left(t-\left(a+K\frac{a}{4C}\right)\right)\leq\frac{a}{4C}$ and $0\leq\left(t-\left(a+2\frac{a}{4C}\right)\right)$. Therefore there exists $k\in\{2,\ldots,K\}$ so that
$\delta:=t-\left(a+k\frac{a}{4C}\right)\quad\textrm{ fulfills }\quad 0\leq\delta\leq\frac{a}{4C},$
since one will certainly land in $[0,\frac{a}{4C}]$ with step size $\left(\frac{a}{4C}\right)$. The asseration follows from Lemma 3. ∎
Remark.
Note that
$a+K\frac{a}{4C}\geq b+\frac{a}{4C}$
holds for $C=1$ with $K\geq\lceil\frac{4b}{a}\rceil-3$. For $C=(n-1)$ we need $K\geq\lceil\frac{4(n-1)b}{a}\rceil-4n-3$.
Theorem 3.
Let $m,n\in\mathbb{N}$, $n,m\geq 2$ and $F\colon\mathbb{R}^{n}_{+}\to\mathbb{S}_{m}$ be continously differentiable, convex and monotonically non-increasing. If
$\mathcal{F}^{\prime}\left(z_{j,k}\right)d_{j}\npreceq 0\quad\textrm{for all}\quad k\in\{2,\ldots,K\},\leavevmode\nobreak\ j\in\{1,\ldots,n\},$
where
$z_{j,k}:=\frac{a}{2}e_{j}^{\prime}+\left(a+k\frac{a}{4}\right)e_{j}\in\mathbb{R}^{n}_{+},\leavevmode\nobreak\ d_{j}:=\frac{2b-a}{a}e_{j}^{\prime}-\frac{1}{2}e_{j}\in\mathbb{R}^{n}$
and $K:=\max\left(\lceil\frac{4b}{a}\rceil-3,2\right)$, then the following holds:
  • (a)
    $F(x)^{\prime}\in\mathcal{L}\left(\mathbb{R}^{n},\mathbb{S}_{m}\right)$ is injective for all $x\in[a,b]^{n}$ and for all $0\neq d\in\mathbb{R}^{n}$,
    $\frac{\|F(x)^{\prime}d\|_{2}}{\|d\|_{\infty}}\geq\lambda:=\min_{j=1,\ldots,n\atop k=2,\ldots,K}\lambda_{\max}\left(F^{\prime}(z_{j,k})d_{j}\right)>0.$
  • (b)
    $F\colon\mathbb{R}^{n}_{+}\to\mathbb{S}^{m}$ is injective and
    $\|F(x_{1})-F(x_{2})\|_{2}\geq\lambda\|x_{1}-x_{2}\|_{\infty}$
    for all $x_{1},x_{2}\in[a,b]^{n}$ i.e. the inverse problem
    determine $\textrm{determine }\hat{x}\in[a,b]^{n}\textrm{ from the knowledge of }F(\hat{x}).$
    is uniquely solvable.
Proof.
Follows directly from Lemma 2 and Corollary 1. ∎
This completes the proof of Theorem 1, since the constructed forward map $\mathcal{F}$ is continously differentiable, convex and monotonically non-increasing by Lemma 1. Note that if $\mathcal{F}$ fulfills the assumptions of Theorem 2, then $\mathcal{F}$ is already injective, i.e. the inverse problem of reconstructing ${\hat{\gamma}}$ from $\mathcal{F}({\hat{\gamma}})$ is uniquely solvable, since the assumptions of Theorem 2 implie the ones of Theorem 1.
3 Numerical results
In the previous section, we present a criterion to ensure that a given number of electrodes is sufficient for the inverse problem to be uniquely solvable, and reformulate the reconstruction problem as a convex semidefinite optimisation problem.
Figure 3: Resolution dimension $n=3$ and $m=5$ electrodes.
We have managed to overcome the non-convexity of the natural data fitting approach, that is illustrated in Figure 2. For a given resolution and electrodes with fixed sizes and positions one can check 4 for unique solvability and 5 for the convex reformulation. However, there is no guarantee that the criteria will be satisfied with finitely many electrodes. It remains to check the criteria 4 and 5
$\mathcal{F}^{\prime}\left(z_{j,k}\right)d_{j}\npreceq 0\quad\textrm{for all}\quad k\in\{2,\ldots,K\},\leavevmode\nobreak\ j\in\{1,\ldots,n\}$
numerically with $z_{j,k}$ and $d_{j}$ chosen accordingly for all $k\in\{2,\ldots,K\},\leavevmode\nobreak\ j\in\{1,\ldots,n\}$ as in Theorem 1 and Theorem 2. Here we consider the simple example of a small circle in a large circle. We divide $\Gamma=\Gamma_{1}\cup\ldots\cup\Gamma_{n}$ equidistant and place electrodes $\mathcal{E}_{1},\ldots,\mathcal{E}_{m}$ uniformly at the boundary as in Figure 3.
3.1 Condition for uniqueness and Lipschitz stability
In this section, we check for different resolutions whether there is a sufficient number of electrodes such that condition 4 is satisfied for the geometry of a small circle inside a large circle. We compute
$\lambda=\min_{j=1,\ldots,n\atop k=2,\ldots,K}\lambda_{\max}\left(\mathcal{F}^{\prime}(z_{j,k})d_{j}\right)$
with $z_{j,k},d_{j}$ for $k\in\{2,\ldots,K\},\leavevmode\nobreak\ j\in\{1,\ldots,n\}$ chosen accordingly to Theorem 1 and check if $\lambda\in\mathbb{R}$ is positive. In fact, one can observe that for a fixed resolution with a growing number of electrodes, the stability constant $\lambda$ is monotonically growing as one can see in Figure 4.
Figure 4: The stability constant $\lambda$ with a fixed resolution dimension $n=3$ for increasing number of electrodes.
In Figure 5, we depict the smallest number electrodes so that $\lambda$ is positive and in particular criterion 4 is satisfied. We therefore achieve unique reconstruction of the corrosion parameter and Lipschitz stability, i.e.
$\|\mathcal{F}(\gamma_{1})-\mathcal{F}(\gamma_{2})\|_{2}\geq\lambda\|\gamma_{1}-\gamma_{2}\|_{\infty}.$
Refer to captionRefer to caption
Figure 5: Plot of the required number of electrodes for criterion 4 and the stability constant $\lambda$ for different resolutions.
However, the inverse stability constant $1/\lambda$ goes to infinity for higher resolutions as shown in Figure 5. Consequently, the ill-posedness of the inverse problem remains a major challenge. In Figure 4, one can observe that $\lambda$ stays substantially small even if one increases the number of electrodes for a fixed resolution.
3.2 Condition for convex reformulation
Uniqueness and Lipschitz stability do not ensure the existence of a global convergent reconstruction algorithm. For the convex semidefinite reformulation, we need to check condition 5 from Theorem 2 with the tighter assumptions. In Figure 6, we depict, for a given resolution, the required number of electrodes such that criterion 5 is fulfilled. We achieve a convex reformulation
$\texttt{minimize}\leavevmode\nobreak\ \sum_{i=1}^{n}|\gamma_{i}|\leavevmode\nobreak\ \textrm{subject to}\leavevmode\nobreak\ \gamma\in[a,b]^{n},\leavevmode\nobreak\ \mathcal{F}(\gamma)\preceq\mathcal{F}(\hat{\gamma}).$ (9)
of the inverse problem 3.
Refer to captionRefer to caption
Figure 6: Plot of the required number of electrodes for criterion 5 and the stability constant $\lambda$ for different resolutions.
However, the stability constant $\lambda$ decreases significantly faster in this setting. This is to be expected since in Theorem 2 one needs to test the criterion at more evaluation points.
Figure 7: Plot of admissible set without and with noise.
In Figure 7, we take a look at the admissible set
$\{\gamma\in[80,120]^{2}\leavevmode\nobreak\ :\leavevmode\nobreak\ \mathcal{F}(\gamma)\preceq\mathcal{F}(\hat{\gamma})\}$
for the convex problem 9 with $\hat{\gamma}=(100,100)$, resolution dimension $n=2$ and $m=5$ tested electrodes. We also add a random noise to $\mathcal{F}(\hat{\gamma})$ s.t. $\|Y^{\delta}-\mathcal{F}(\hat{\gamma})\|_{2}\leq\delta$ and compute
$\{\gamma\in[80,100]^{2}\leavevmode\nobreak\ :\leavevmode\nobreak\ \mathcal{F}(\gamma)\preceq Y^{\delta}+\delta I\}$
with $\delta=0,10^{-5},10^{-4},10^{-3}$.
4 Conclusions
This work presents an explicit criterion for the required number of electrodes for unique solvability and a convex reformulation of an elliptic inverse coefficient problem. Once a suitable geometry has been chosen and electrodes of fixed sizes and positions are placed at the boundary, then the forward problem can be simulated. By performing a finite number of forward evaluations, it is possible to verify the conditions 4 and 5 by numerically calculating a value $\lambda\in\mathbb{R}$ and checking if it is positive. Ultimately, the problem can be rewritten as a uniquely solvable, convex, semidefinite program, where the stability constant of the inverse problem turns out to be $1/\lambda$. In the case of a breach of the conditions 4 and 5, i.e. if $\lambda\leq 0$, additional electrodes may be applied and the procedure repeated until $\lambda>0$. The results of the experiments in Figure 4, Figure 5 and Figure 6 indicate that the criteria are satisfied for a sufficient number of electrodes.